Tik-61.181: Bioinformatics Kaski, Mannila, Nikkilä 8th November 2000

Some preliminary topics of the project work

1) Your own topic (creatively utilized opportunity to invent your own project work is positively taken into account when the grades are decided)

– the course material(books and articles) form an excellent base for project work

– personnel of the course has some data (at least expression data) and sequences and other biological data as well as some program packages can quite easily be found from the WWW (assistant will shortly provide some links to these sites)

2) Analysis of the expressions of yeasts known genes in different treatments (Visual-Gene, for group of two people)

– data includes all the known genes of the yeast Saccharomyces Cerevisiae and their expression profiles in several treatments

– the program runs on Windows environment and the license includes 50 start ups of the program

3) Analysis of the expressions of the yeasts all genes in diauxic shift (VisualGene, for group of two people)

– data includes the all genes of same yeast, but only one treatment: diauxic shift

– in addition the data has variables concerning the micro array technique

– the program runs on Windows environment and the license includes 50 start ups of the program

4) Gene expression data analysis with SOMToolbox

– data to be decided